OpenStudy (anonymous):

How do I convert 1.2 x 10^-5kg into ng?

OpenStudy (saifoo.khan):

ng or mg?

OpenStudy (anonymous):

ng

OpenStudy (dumbcow):

nice...nanograms

OpenStudy (anonymous):

nano nano

OpenStudy (dumbcow):

i believe there are 1,000,000,000 ng in 1 gram and there are 1000g in 1 kg

OpenStudy (anonymous):

my god! if that's the case I'll never figure out this problem!

OpenStudy (dumbcow):

not at all...just count the zeroes (1.2X10^-5)*(1x10^9)*(1X10^3)

OpenStudy (anonymous):

I am still totally lost...

OpenStudy (dumbcow):

ok i just converted the numbers into scientific notation a ng is much much smaller than a kg so we multiply to make our number bigger (1.2X10^-5)kg * (1X10^3)g/kg * (1x10^9)ng/g = 1.2X10^7 ng

OpenStudy (anonymous):

ohh okay!

OpenStudy (dumbcow):

i added the exponents to get the 7: -5+9+3 = -5+12 = 7

OpenStudy (anonymous):

Oh I understand

OpenStudy (dumbcow):

:)